OpenStudy (anonymous):

If a, b are the solutions of the equation log base of 2(5 − 2x) + log base of 2(5 + 2x) = 2 , find the value of the product ab. I know the law is log base 2 (5-2x)(5+2x)=2 So After I foil, I get 25+10x-10x-4x^2. I then set -4x^2+25=2 and solve for x. I pretty much get x=square root of 23/4 however that is not an answer choice. Did I do something wrong?

OpenStudy (campbell_st):

well what you forgot was to raise the right hand side of the equation as a power of 2 so it should be -4x + 25 = 2^2 -4x + 25 = 4 now solve for x

OpenStudy (anonymous):

Thank you so much @campbell_st

OpenStudy (campbell_st):

hope it helps... remember for equations its the same thing on both sides of the =